compute the electric force between two charges of 5×10^-9 C and -3×10-8 C which are reparated by r=10 cm


1
Expert's answer
2022-02-28T13:40:38-0500

The Coulomb's law says

"F=k\frac{q_1q_2}{r^2}\\
=9*10^9*\frac{5*10^{-9}*3*10^{-8}}{0.1^2}=1.35*10^{-4}\;\rm N"